Blocking refers to the idea that the variability in a variable can be reduced by segmenting the data by some other variable. The data in the accompanying table represent the recumbent length​ (in centimeters) of a sample of 10 males and 10 females who are 40 months of age. Complete parts​ (a) through​ (d).

Blocking refers to the idea that the variability in a variable can be reduced by segmenting the data by some other variable. The data in the accompanying table represent the recumbent length​ (in centimeters) of a sample of 10 males and 10 females who are 40 months of age. Complete parts​ (a) through​ (d).

Males   Females
104.0 102.7
93.7   100.4
98.3   102.8
86.2   98.1
90.7   95.4
94.4   100.8
97.6   96.3
100.6   105.0
103.0 106.5
100.9   114.5

​(a) Determine the standard deviation of recumbent length for all 20 observations.

cm ​(Round to two decimal places as​ needed.)

​(b) Determine the standard deviation of recumbent length for the males.

.......cm ​(Round to two decimal places as​ needed.)

​(c) Determine the standard deviation of recumbent length for the females.

...........cm ​(Round to two decimal places as​ needed.)

​(d) What effect does blocking by gender have on the standard deviation of recumbent length for each​ gender?

...........cm ​(Round to two decimal places as​ needed.)

2, Suppose that a customer is purchasing a car. He conducts an experiment in which he puts 10 gallons of gas in the car and drives it until it runs out of gas. He conducts this experiment 15 times on each car and records the number of miles driven.

Car 1 Car 2
220 239
212 246
226 219
215 238
248 251
273 155
261 230
258 274
255 154
243 291
250 273
247 319
298 311
255 291
293 285

Median for Car 1

M = mi​ / 10 gal

​(Type an integer or decimal rounded to one decimal place as​needed.)

Median for Car 2

Me =   mi​ / 10 gal

​(Type an integer or decimal rounded to one decimal place as​needed.)
Answer #1

Q1

Formula - SD(\sigma ) = \sqrt[2]({(\sum (X- \bar{X})^{2})}/(N-1))

where,

SD is standard deviation

X is sample data given

Xbar is Mean of sample data

N is Number of sample data

1(a) - Standard Deviation for all 20 observations - Answer is 6.12 (Solution attached)

ALL 20 OBSERVATIONS
S No X X - Xbar (X-Xbar)2
1 104 4.405 19.40403
2 93.7 -5.895 34.75103
3 98.3 -1.295 1.677025
4 86.2 -13.395 179.426
5 90.7 -8.895 79.12102
6 94.4 -5.195 26.98802
7 97.6 -1.995 3.980025
8 100.6 1.005 1.010025
9 103 3.405 11.59403
10 100.9 1.305 1.703025
11 102.7 3.105 9.641025
12 100.4 0.805 0.648025
13 102.8 3.205 10.27203
14 98.1 -1.495 2.235025
15 95.4 -4.195 17.59802
16 100.8 1.205 1.452025
17 96.3 -3.295 10.85703
18 105 5.405 29.21403
19 106.5 6.905 47.67903
20 114.5 14.905 222.159
Xbar(Mean) 99.595
∑(X-Xbar)2 711.4095
S.D. (as per formula) SQRT(711.4095/19) 6.12

1(b) - Standard Deviation for Males observations - Answer is 5.67 (Solution Attached)

MALES
S No X X - Xbar (X-Xbar)2
1 104 7.06 49.8436
2 93.7 -3.24 10.4976
3 98.3 1.36 1.8496
4 86.2 -10.74 115.3476
5 90.7 -6.24 38.9376
6 94.4 -2.54 6.4516
7 97.6 0.66 0.4356
8 100.6 3.66 13.3956
9 103 6.06 36.7236
10 100.9 3.96 15.6816
Xbar(Mean) 96.94
∑(X-Xbar)2 289.164
S.D. (as per formula) SQRT(289.164/9) 5.67

1(c) - Standard Deviation for Females observations - Answer is 5.59 (Solution Attached)

FEMALES
S No X X - Xbar (X-Xbar)2
1 102.7 0.45 0.2025
2 100.4 -1.85 3.4225
3 102.8 0.55 0.3025
4 98.1 -4.15 17.2225
5 95.4 -6.85 46.9225
6 100.8 -1.45 2.1025
7 96.3 -5.95 35.4025
8 105 2.75 7.5625
9 106.5 4.25 18.0625
10 114.5 12.25 150.0625
Xbar(Mean) 102.25
∑(X-Xbar)2 281.265
S.D. (as per formula) SQRT(281.265/9) 5.59

1(d) - Answer is - The standard deviation is lower for each individual group than it is for the genders combined.

Q2

For median, first step is always to sort given sample data in ascending order

CAR 1 Ascending CAR 2 Ascending
220 212 239 154
212 215 246 155
226 220 219 219
215 226 238 230
248 243 251 238
273 247 155 239
261 248 230 246
258 250 274 251
255 255 154 273
243 255 291 274
250 258 273 285
247 261 319 291
298 273 311 291
255 293 291 311
293 298 285 319

Formula for Median - when n (sample size) is odd number.

Median = (n+1)/2 th term

Formula for Median - when n (sample size) is even number.

Median: [(n/2) th term + {(n/2)+1} th term ] / 2

In our case, n is 15 which is odd number, ​

So,

Median for Car 1 = (15+1)/2 th term i.e. 8th term.

Median for Car 1 = 250

Median for Car 2 = (15+1)/2 th term i.e. 8th term.

Median for Car 2 = 251